OpenStudy (angel110):

1. If the height h of a triangle is 3 inches less than the length of the base b, and the area A of the triangle is 19 times the length of the base, find b and h.

OpenStudy (jdoe0001):

h = Height b = Base A = Area -------------------------------- we know that the Height is 3" less than Base h = b -3 we also know that the Area of the triangle is 19 times Base \(Area = \cfrac{1}{2}\times base \times height \implies 19base\implies 19b\) so we can say that \(\bf 19b = \cfrac{1}{2}bh\\ \textit{but we know that } h = b-3\\ 19b = \cfrac{1}{2}bh\implies 19b = \cfrac{1}{2}b(b-3) \implies 19b = \cfrac{b(b-3)}{2}\\ \textit{let's multiply both sides by 2}\\ 19b = \cfrac{b(b-3)}{2} \implies 38b = b^2-3b\) solve for "b" to find the Base keep in mind that Height is just Base -3